# Entanglement measures for detectability

### v2

@@ C.3 Proof of Theorem 9
$\Omega_{\alpha,{\cal D}}[\rho_{AC}]:=\mathop{\rm sgn}(\alpha-1)\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (149)
@@
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
@@
$\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}(\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*})\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
@@
Since $Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})$ and $(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})$ are density matrices, we have
@@
$=$ $-\operatorname{\mathrm{Tr}}_{AC}\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $-\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Bigg{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
@@
$=$ $(\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})+\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}(\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1)$
@@
$=$ $(\operatorname{\mathrm{Tr}}Z^{\prime}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})+\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha}(\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1).$ (155)
@@ D.1 Proof of (115)
for a general state on ${\mathcal{H}}^{\otimes n}$. We choose a basis $\{|u_{\lambda,j}\rangle\}$ of ${\cal U}_{\lambda}$ such that $|u_{\lambda,j}\rangle$ is a eigenvector of $\rho_{\lambda}$. We define the pinching map $\Gamma_{\rho,n}$ as
@@
for a general state $\rho^{\prime}$ on ${\mathcal{H}}^{\otimes n}$. Due to the above mentioned structure, states $\Gamma_{\rho,n}(\sigma^{\otimes n})$ and $\rho^{\otimes n}$ are commutative with each other. We denote the dimension of $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ by $d_{n}$. Since $\dim{\cal U}_{\lambda}\leq(n+1)^{d(d-1)/2}$ and $|Y_{d}^{n}|\leq(n+1)^{d-1}$ [51, (6.16) and (6.18)], we have
@@
Since it is natural to define $(I-P_{1})\cap(I-P_{2}):=(I-P_{1})(I-P_{2})$, the above definition for $P_{1}\cup P_{2}$ is also natural. When $\{|u_{j}\rangle\}_{j}$ is a common eigenvector system of $P_{1}$ and $P_{2}$, they can be written as $P_{k}=\sum_{j\in J_{k}}|u_{j}\rangle\langle u_{j}|$. Then, we have $P_{1}\cup P_{2}=\sum_{j\in J_{1}\cup J_{2}}|u_{j}\rangle\langle u_{j}|$. Based in this definition, we define the operator $T_{n,\alpha,r}$ as
@@
Although $T_{n}(\sigma,R_{n,\alpha,r})$ is an operator on ${\mathcal{H}}^{\otimes n}$, it can be written by using a projection operator $\oplus_{\lambda\in Y_{d}^{n}}P_{\sigma,n,\lambda}$ on $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ as
@@
Hence, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}(\bigcup_{\sigma\in S}P_{\sigma,n,\lambda})\otimes P({\cal V}_{\lambda}).$ (166)
Since all $P_{\sigma,n,\lambda}$ are commutative with each other, we can choose $\dim{\cal U}_{\lambda}$ elements $\sigma_{\lambda,j}$ in $S$ such that
$\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}=\bigcup_{j}P_{\sigma_{\lambda,j},n,\lambda}.$ (167)
@@
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}(\bigcup_{j}P_{\sigma_{\lambda,j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j}\bigoplus_{\lambda\in Y_{d}^{n}}(P_{\sigma_{\lambda^{\prime},j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r}).$ (168)
@@
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\leq\operatorname{\mathrm{Tr}}\rho^{\otimes n}\sum_{\lambda^{\prime}\in Y_{d}^{n},j}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\leq$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j}\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\leq$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$\leq$ $d_{n}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$=$ $d_{n}\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}\big{)}$
@@
$\leq$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n}(\sigma,R_{n,\alpha,r}))$
$\leq$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}$
@@
Thus, we have
@@
$\leq$ $\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$ (172)
Thus,
@@
$\cdot\big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}\big{)}$
@@
$\cdot\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}))^{\alpha-1}$
$\leq$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
@@
$=$ $d_{n}^{1-\alpha}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n}.$ (173)
@@
$=$ $d_{n}^{2}\exp\Big{(}n\frac{(1-\alpha)E_{\alpha}^{*}(\rho)-\alpha r}{1-\alpha}\Big{)}.$ (174)
Since arbitrary $\alpha\in[0,1]$ satisfies (171), (174) yields
@@
D.2 Proof of (116)
Here, we employ the idea invented in [58], whose detail is explained in [42, Section 3.8]. We choose a test $T_{n}$ such that

### v3

@@ C-C Proof of Theorem 12
$\Omega_{\alpha,{\cal D}}[\rho_{AC}]$
::=:= $\mathop{\rm sgn}(\alpha-1)\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}.$ (157)
@@
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}_{AC}\rho_{AC}\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
@@
$\otimes\operatorname{\mathrm{Tr}}_{B}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $\mathop{\rm sgn}(\alpha-1)\operatorname{\mathrm{Tr}}(\sum_{c\in{\cal C}}X_{A|c}\otimes\rho_{B|c}^{*})$
$\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
@@
Since $Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}$ and $(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})$ are density matrices, we have
@@
$=$ $-\operatorname{\mathrm{Tr}}_{AC}\sigma_{AC}[\{X_{A|c}^{*}\}_{c}]\sum_{c\in{\cal C}}|c\rangle\langle c|\otimes$
$\Big{(}\operatorname{\mathrm{Tr}}_{B}\Big{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$-\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c^{\prime}\in{\cal C}}X_{A|c^{\prime}}\otimes\rho_{B|c^{\prime}}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}$
$(I_{A}\otimes\rho_{B|c}^{*})\Big{)}$
$=$ $-\operatorname{\mathrm{Tr}}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}$
$\Bigg{(}\Big{(}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{(}\sum_{c\in{\cal C}}X_{A|c}^{*}\otimes\rho_{B|c}^{*}\Big{)}\rho_{AB}^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}$
@@
$=$ $(\operatorname{\mathrm{Tr}}Z_{*}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})$
$+\operatorname{\mathrm{Tr}}Z_{*}^{\alpha}(\operatorname{\mathrm{Tr}}(Z_{*}^{\alpha}/\operatorname{\mathrm{Tr}}Z_{*}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1)$
@@
$=$ $(\operatorname{\mathrm{Tr}}Z^{\prime}Z^{\alpha-1})(\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{\alpha-1}{\alpha}}((\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1-\alpha}{\alpha}}-(\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{1-\alpha}{\alpha}})$ (163)
$+\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha}(\operatorname{\mathrm{Tr}}({Z^{\prime}}^{\alpha}/\operatorname{\mathrm{Tr}}{Z^{\prime}}^{\alpha})^{\frac{1}{\alpha}}(Z^{\alpha}/\operatorname{\mathrm{Tr}}Z^{\alpha})^{\frac{\alpha-1}{\alpha}}-1).$ (164)
@@ D-A Proof of (123)
for a general state on ${\mathcal{H}}^{\otimes n}$, where $P({\cal K})$ expresses the projection to the subspace ${\cal K}$. We choose a basis $\{|u_{\lambda,j}\rangle\}_{j=1}^{d_{\lambda}}$ of ${\cal U}_{\lambda}$ such that $|u_{\lambda,j}\rangle$ is an eigenvector of $\rho_{\lambda}$, where $d_{\lambda}:=\dim{\cal U}_{\lambda}$. We define the pinching map $\Gamma_{\rho,n}$ as
@@
for a general state $\rho^{\prime}$ on ${\mathcal{H}}^{\otimes n}$. Due to the above mentioned structure, states $\Gamma_{\rho,n}(\sigma^{\otimes n})$ and $\rho^{\otimes n}$ are commutative with each other. We denote the dimension of $\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}$ by $d_{n}$. Since $\dim{\cal U}_{\lambda}\leq(n+1)^{d(d-1)/2}$ and $|Y_{d}^{n}|\leq(n+1)^{d-1}$ [70, (6.16) and (6.18)], we have
@@
Since it is natural to define $(I-P_{1})\cap(I-P_{2}):=(I-P_{1})(I-P_{2})$, the above definition for $P_{1}\cup P_{2}$ is also natural. When $\{|u_{j}\rangle\}_{j}$ is a common eigenvector system of $P_{1}$ and $P_{2}$, they can be written as $P_{k}=\sum_{j\in J_{k}}|u_{j}\rangle\langle u_{j}|$. Then, we have $P_{1}\cup P_{2}=\sum_{j\in J_{1}\cup J_{2}}|u_{j}\rangle\langle u_{j}|$. Based in the definition (172), we define the operator $T_{n,\alpha,r}$ as
@@
Although $T_{n}(\sigma,R_{n,\alpha,r})$ is an operator on ${\mathcal{H}}^{\otimes n}$, it can be written by using a projection operator $P_{\sigma,n,\lambda}$ on ${\cal U}_{\lambda}$ as
@@
That is, we choose the projection operator $P_{\sigma,n,\lambda}$ as the above. Hence, we have
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}\Big{(}\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}\Big{)}\otimes P({\cal V}_{\lambda}).$ (175)
Since all projections $P_{\sigma,n,\lambda}$ are diagonalized with the basis $\{|u_{\lambda,j}\rangle\}_{j=1}^{d_{\lambda}}$ of ${\cal U}_{\lambda}$, it is sufficient to clarify whether the range of $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$. If the range of $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$, there exists an element $\sigma_{\lambda,j}\in S$ such that $P_{\sigma_{\lambda,j},n,\lambda}$ contains $|u_{\lambda,j}\rangle$. If the range of $P_{\sigma,n,\lambda}$ does not contain $|u_{\lambda,j}\rangle$, no element $\sigma\in S$ satisfy the condition $P_{\sigma,n,\lambda}$ contains $|u_{\lambda,j}\rangle$. Therefore, we can choose $d_{\lambda}$ elements $\sigma_{\lambda,1},\ldots,\sigma_{\lambda,d_{\lambda}}$ in $S$ such that
$\bigcup_{\sigma\in S}P_{\sigma,n,\lambda}=\bigcup_{j\in[d_{\lambda}]}P_{\sigma_{\lambda,j},n,\lambda},$ (176)
where $[d_{\lambda}]:=\{1,\ldots,d_{\lambda}\}$. Here, the elements $\sigma_{\lambda,j}$ are not necessarily distinct, in general.
@@
$T_{n,\alpha,r}=\bigoplus_{\lambda\in Y_{d}^{n}}\Big{(}\bigcup_{j\in[d_{\lambda}]}P_{\sigma_{\lambda,j},n,\lambda}\Big{)}\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\bigoplus_{\lambda\in Y_{d}^{n}}(P_{\sigma_{\lambda^{\prime},j},n,\lambda})\otimes P({\cal V}_{\lambda})$
$=$ $\bigcup_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r}).$ (177)
@@
$\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n,\alpha,r}\stackrel{{\scriptstyle(a)}}{{\leq}}\operatorname{\mathrm{Tr}}\rho^{\otimes n}\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$=$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\operatorname{\mathrm{Tr}}\rho^{\otimes n}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\sum_{\lambda^{\prime}\in Y_{d}^{n},j\in[d_{\lambda}]}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
=$\stackrel{{\scriptstyle(d)}}{{=}}$ $d_{n}\big{(}\max_{\sigma\in S}\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}\big{)}$
@@
Each step can be shown as follows. Step $(a)$ follows from (177). Step $(b)$ can be shown as follows. Since the condition $\Gamma_{\rho,n}(\sigma^{\otimes n})\geq e^{nR}\rho^{\otimes n}$ implies $\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}\geq e^{\alpha nR}(\rho^{\otimes n})^{\alpha}$, the definition (171) guarantees that the inequality
$T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})\rho^{\otimes n}$
$=$ $T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})(\rho^{\otimes n})^{1-\alpha}(\rho^{\otimes n})^{\alpha}$
$\leq$ $\operatorname{\mathrm{Tr}}T_{n}(\sigma_{\lambda^{\prime},j},R_{n,\alpha,r})\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}$
$=$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma_{\lambda^{\prime},j}^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{-n\alpha R_{n,\alpha,r}}.$
Step $(c)$ follows from the relation $\sum_{\lambda\in Y_{d}^{n}}d_{\lambda}=\dim\oplus_{\lambda\in Y_{d}^{n}}{\cal U}_{\lambda}=d_{n}$. Step $(d)$ follows from the definition of $R_{n,\alpha,r}$.
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})(I-T_{n}(\sigma,R_{n,\alpha,r}))$
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $\operatorname{\mathrm{Tr}}\Gamma_{\rho,n}(\sigma^{\otimes n})^{\alpha}(\rho^{\otimes n})^{1-\alpha}e^{n(1-\alpha)R_{n,\alpha,r}}$
@@
Each step can be shown as follows. Step $(a)$ follows from the relation $T_{n}(\sigma,R_{n,\alpha,r})\leq T_{n,\alpha,r}$, which is shown by the relation (175). Step $(b)$ follows from (171) in the same way as Step $(b)$ of (178). Thus, we have
@@
$\leq$ $\Big{(}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}^{\alpha-1}.$ (181)
Thus, we have
@@
$\cdot\Big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
@@
$\cdot\Big{(}\Gamma_{\rho,n}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})\Big{)}^{\alpha-1}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n}^{1-\alpha}\operatorname{\mathrm{Tr}}((\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\sigma^{\otimes n}(\rho^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
@@
$=$ $d_{n}^{1-\alpha}(\operatorname{\mathrm{Tr}}(\rho^{\frac{1-\alpha}{2\alpha}}\sigma\rho^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (182)
where Step $(a)$ follows from (181).
@@
$=$ $d_{n}^{2}\exp\Big{(}n\frac{-(1-\alpha)E_{\alpha}^{*}(\rho)+\alpha r}{1-\alpha}\Big{)}.$ (183)
Since an arbitrary number $\alpha\in[0,1]$ satisfies (180), (183) yields
@@
Remark 4
Although the above presented proof employs several techniques that come from prior works, this proof is not simple combination of techniques that come from prior works. Our test is given in (175), and it has not been introduced in existing references [21, 22]. In fact, the method by [21, 22] cannot be applied to the case when $H_{0}$ hypothesis is the set of all separable states over $n$ bipartite systems, as opposed to an IID $n$-fold tensor product of a separable state. To extend quantum Sanov theorem to this case, our obtained evaluation is essentially needed. To upper bound the error probability, we have invented a bounding method different from the existing bounding method for the simple hypothesis [78] because the existing bounding method does not work in this case and it is related to Petz relative Rényi entropy. In fact, the above evaluation requires new careful handling the parameter $\alpha$ due to the matrix convexity.
D-B Proof of (124)
Here, we employ the idea invented in [77], whose detail is explained in [12, Section 3.8]. We choose a test $T_{n}$ such that